OpenStudy (candy13106):

use the quadratic formula to solve the equation. round your solution to the nearest hundredth x^2+6x-10=0

OpenStudy (candy13106):

can u help me plz @Here_to_Help15

OpenStudy (candy13106):

idk

OpenStudy (candy13106):

ok idk how to do that i learned it the other way

OpenStudy (candy13106):

ok imk so sorry

OpenStudy (johnweldon1993):

Okay, so quadratic formula lol \[\large \frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\]

OpenStudy (candy13106):

watever idk need to do my homework

OpenStudy (johnweldon1993):

The equation x^2 + 6x - 10 = 0 We have a = 1 b = 6 c = -10 So plug all those into the quadratic equation \[\large \frac{-6 \pm \sqrt{6^2 - 4(1)(-10)}}{2(1)}\]

OpenStudy (candy13106):

ok

OpenStudy (johnweldon1993):

Simplify it a little bit \[\large \frac{-6 \pm \sqrt{36 - (-40)}}{2}\]

OpenStudy (candy13106):

ok

OpenStudy (johnweldon1993):

\[\large \frac{-6 \pm \sqrt{76}}{2}\] Soo now...what is the square root of 76?

OpenStudy (candy13106):

8.7

OpenStudy (johnweldon1993):

Okay, so we have \[\large \frac{-6 \pm 8.717797887081348}{2}\] *sorry just wrote it out lol Now break that into 2 equations \(\large \frac{-6 + 8.717797887081348}{2}\) and \(\large \frac{-6 - 8.717797887081348}{2}\) solve those both and you're good :)

OpenStudy (candy13106):

1.35 and -7.35 right?

OpenStudy (johnweldon1993):

Perfect :)

OpenStudy (candy13106):

that is right

OpenStudy (johnweldon1993):

I checked the first one and it was right so I'll assume the second is as well :D

OpenStudy (candy13106):

ok thanks